The annual wages X of the staffs of collage of natural sciences are normally


distributed with mean µ and standard deviation σ. Knowing that P(X > 42320) =


0.0853 and P(X < 41230) = 0.6103, find µ and σ.

1
Expert's answer
2022-04-12T17:24:51-0400

"P(X>42320)=0.0853;\\
P(X<42320)=P(Z<z_1)=\\
=1-0.0853=0.9147;\\
P(X<41230)=0.6103;\\
z_1=\cfrac{42320-\mu}{\sigma};\\
z_2=\cfrac{41230-\mu}{\sigma}."

From z-table we find "P(1.37)=0.9147,P(0.28)=0.6103."

So,

"\begin{cases}
 \cfrac{42320-\mu}{\sigma}=1.37\\
\cfrac{41230-\mu}{\sigma}=0.28
\end{cases}"

"\begin{cases}
42320-\mu=1.37\sigma\\
41230-\mu=0.28\sigma
\end{cases}"

"\begin{cases}
\mu=42320-1.37\sigma\\
\mu=41230-0.28\sigma
\end{cases}"

"42320-1.37\sigma=41230-0.28\sigma\\
1.09\sigma=1090\\
\sigma=1000\\
\mu=41230-0.28\cdot1000=40950."